OpenStudy (jnstalex):

Simplify 2 to the 5th over 3 squared all raised to the 4th power.

OpenStudy (stickman1998):

\[((2^5)/3^2)^4\]

OpenStudy (stickman1998):

Like this?

OpenStudy (dopeness_):

Yes, that's the correct format @stickman1998 .

OpenStudy (dopeness_):

@Jnstalex What's \[2\times2\times2\times2\times2=?\]

TheSmartOne (thesmartone):

Hint: \(\Large\bf( \frac{a}{b})^c = \frac {a^c}{b^c}\) \(\Large\bf (x^y)^z = x^{yz}\)